-3x+2y=-6
-2x+y=6

Answer by stanbon(75887)

-3x+2y=-6
-2x+y = 6
===========
Multiply thru the bottom equation by 2:
-3x + 2y = -6
-4x + 2y = 12
-------------------
Subtract and solve for "x":
x = -18
------
Solve for "y" using -2x+y = 6
y = 2x + 6
---
y = 2(-18)+6
y = -30
============
Cheers,
Stan H.
============